5+4a=5a+2
solve eqautions with variables on both sides

Answer by tutorcecilia(2152)

5+4a=5a+2 Isolate the variable to one side of the equation
5-5+4a=5a+2-5 Subtract, add, multiply or divide to isolate the variable
4a=5a-3
4a-5a=5a-5a-3
-a=-3
-1a/-1=-3/-1
a=3
.
Check your answer by plugging a =3 back into the original equation
.
5+4(3)=5(3)+2
5+12 = 15+2
17=17 Checks out that a = 3